OpenStudy (anonymous):

Frank parachutes from a plane at a height of 25,000 ft above sea level. Several minutes later he notes from an instrument on his wrist that he had dropped 15,000 ft. How far above sea level is he now?

OpenStudy (anonymous):

10,000

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can you help me with this? Which equation models the problem below if c represents the number of puppies whose eyes are closed? There are 15 puppies in a photograph, and 11 of them have their eyes open. How many puppies have their eyes closed?

OpenStudy (anonymous):

c – 11 = 15 11 – c = 15 15 + c = 11 11 + c = 15

OpenStudy (anonymous):

Are those your choices?

OpenStudy (anonymous):

I think it's the last one.